OpenStudy (anonymous):

Complete each ordered pair so that is is a solution for the equation: 3x-4y=24 (?,6)

jimthompson5910 (jim_thompson5910):

3x-4y=24 3x-4(6)=24 3x-24 = 24 3x = 24+24 3x = 48 x = 48/3 x = 16 So the ordered pair is (16, 6)

OpenStudy (anonymous):

Thankyou!

jimthompson5910 (jim_thompson5910):

np